# Towards a logic of affordances

### v1

@@ 1. Introduction
The affordances of the environment are what it offers the animal, what it provides or furnishes, either for good or ill …I mean by it something that refers to both the environment and the animal in a way that no existing term does. It implies the complementarity of the animal and the environment. [11, p. 119]
@@
  • a dog is fetching a stick on a meadow,
  • a basketball player is dunking a ball on a playground,
  • a robot is serving a meal in a restaurant,
  • a student is solving a test in a classroom,
  • a mathematician is writing a proof on a blackboard in their office,
@@
In this work, we intend to conceptualize affordances as ternary relations between actors, objects, and environments, taking into account their properties. [^1] Some work toward affordances as ternary phenomena (though not necessarily relations in the standard logical sense) was carried out in [16]. The main goal is to construct a framework for building logic (or logics) for reasoning with affordances. Our pre-theoretical definition of affordance is:
@@
An affordance is an action possibility formed by the relationship among an actor, an object, and their environment.
The relevant components of the approach from this work:
  • Actors and objects have various properties, such as: height, weight, color, agility, abilities, disabilities, skills, shape, and so on.
  • Environments also have properties that may be elementary (in some sense), like: flat, bumpy, stifling, hot, wet, rainy; or more complex: adapted for playing basketball, multi-level, equipped with an elevator, dangerous for health, to mention a few.
@@ 2. Information systems
If we think of a property system as describing possible states of an attribute—such as “color” or “language spoken”—we extend it by the definition of an aggregate structure: An attribute system (A-system) [18] is a structure $\mathcal{S}\coloneqq\left\langle U,\Omega,\{V_{a}:a\in\Omega\},f\right\rangle$ where
@@
    $\Omega$ is a set of property labels or attributes, and $V_{a}$ is a set of possible values of $a\in\Omega$,
@@
    $f\colon U\times\Omega\to\bigcup_{a\in\Omega}2^{V_{a}}$ is a choice function, where $f(x,a)\subseteq V_{a}$. Equivalently, we may define $f\colon U\to\prod_{a\in\Omega}2^{V_{a}}$.
So, if $a\in\Omega$ is a property label weight, then $V_{\mathit{weight}}$ may be a set of rational numbers in some interval that can serve as numerical expressions of the weight of an object (e.g., in kilograms or pounds), or any value that makes sense. It could also be some aggregated value such as “low”, “medium”, “high” etc.
We call $\left\langle U,\Omega,\{V_{a}:a\in\Omega\}\right\rangle$ the skeleton of $\mathcal{S}$. The product $U\times\prod_{a\in\Omega}2^{V_{a}}$ collects all possible vectors of value sets that can be associated with some element of $U$. An information function now picks one element from $\prod_{a\in\Omega}2^{V_{a}}$ for each $x\in U$.
Clearly, an attribute system can be regarded as a collection of the property systems. Düntsch et al., [7] introduce relational attribute system as a relational form of attribute systems, and present formalizations of various interpretations of a statement such as $p\in f(x,a)$. These are closely related to databases with incomplete information investigated by Lipski, [12]. A deduction system for reasoning in relation algebras derived from relational attribute systems was presented in [8].
An element $x\in U$ is called deterministic, if $|f(x,a)|\leq 1$ for all $a\in\Omega$ or every projection of the vector attribute to $x$ is either an empty set or a singleton subset of $V_{a}$. The set of all deterministic elements of $\mathcal{S}$ is denoted by $D_{\mathcal{S}}$. The characterization stems from the fact that the role of the choice function is to narrow down the possibilities for the values of $a$ with respect to the object $x$. If $|f(x,a)|=1$, then we know the exact value of the property $a$ for $x$, or if $|f(x,a)|=\emptyset$, then we know that $x$ does not have this property at all. This is why we call a system with $|f(x,a)|\leqslant 1$ deterministic. If $|f(x,a)|\geq 2$ the set has different possibilities of interpretation, see [7].
If $U$ is finite and $U=D_{\mathcal{S}}$, then $\mathcal{S}$ is called an information system (in the sense of Pawlak, [15]). In this case we may simplify the definition of the information function to have codomain $\bigcup_{a\in\Omega}V_{a}$. Vakarelov, [18] calls these systems “ontological information systems”, and provides logical counterparts of P-systems and A-systems based on Scott’s deduction systems. He continues to develop modal logics for similarity relations on P-systems and A-systems and shows completeness with respect to finite models, continued in [19].
@@
The properties we consider do not have to be in any way simple or elementary. We can consider complex features, such as being adapted to playing basketball. Below, we have an information system for playgrounds, each of which is characterized in terms of three complex attributes: being adapted or not (in symbols: 1 or 0) to playing basketball, soccer, and volleyball.
@@
corresponds to the configuration of four environments in which: in $e_{1}$ one diode is active, in $e_{2}$ no diode is active, in $e_{3}$ two diodes are active, and in $e_{4}$ eight. Even in such a relatively simple environment we have
@@ 2.1. Information systems for environments
    We assume that there is one underlying environment (or context) in some pre-theoretical sense, the elements of $E$ are <<parts>> of the environment, and labels in $\Omega_{E}$ are interpreted as features of these parts. In this case, $V^{E}_{p}$ for $p\in\Omega^{E}$ is the set of values for parts of the environment.
@@
    $\Omega^{E}\coloneqq\{\mathrm{state}\}$, one label interpreted as the state of a diode.
@@
Consequently, we have sixteen different information systems based on the skeleton $\left\langle E,\Omega^{E},\{V^{E}_{\mathrm{state}}\}\right\rangle$. Here, there is one environment, and the elements of $E$ are parts of this environment. So this example is a representative of the perspective (E1) above. $\dashv$
@@
corresponds to the configuration of four environments in which: in $e_{1}$ one diode is active, in $e_{2}$ no diode is active, in $e_{3}$ two diodes are active, and in $e_{4}$ eight. Even in such a relatively simple environment we have
@@ 3. Rough sets
Given an information system $\mathcal{I}\coloneqq\left\langle U,\Omega,\{V_{a}\mid a\in\Omega\},f\right\rangle$, for all $x,y\in U$
$x\mathrel{\theta}y\;\mathrel{\mathord{:}\mathord{\Longleftrightarrow}}\;f(x)=f(y),$
that is, $x$ and $y$ have the same value vector. In this case, we say that $x$ is similar to $y$ (or that $x$ and $y$ are similar). Obviously, in the case of the deterministic systems, $x$ and $y$ are similar if they have precisely the same values for all the property labels in $\Omega$. $\dashv$
Clearly, $\theta$ is an equivalence relation on $U$ whose classes we denote by $x]_{\theta}$ or simply by $[x]$ if no confusion can arise. By means of $\theta$, we can make use of a concept of a rough set to incorporate the granularity of the world into our formalism. Given a subset $X$ of $U$ and an element $a\in U$, when can we say that certainly $a\in X$ or certainly $a\not\in X$? Since we know the world only up to the classes of $\theta$ we can say that
@@
$\overline{X}=\{a_{1},a_{3},a_{4},a_{5},a_{7},a_{8},a_{9}\}\quad\text{and}\quad\underline{X}=\{a_{3}\}.$
In consequence, the boundary of $X$ is the set $\{a_{1},a_{4},a_{5},a_{7},a_{8},a_{9}\}$. $\dashv$
@@ 4. Operationalizing affordances
The basic setup of an affordance relation consists of a set $U$ of an agent’s abilities, a set $E$ of features of the environment, and a binary relation $R\subseteq U\times E$. Chemero, [1, p. 189] writes
@@
$\mathcal{I}_{A}=\left\langle A,\Omega_{A},\{V_{q}^{A}:q\in\Omega_{A}\},f_{A}:A\to\prod_{q\in\Omega_{A}}V_{q}^{A}\right\rangle,\\[5.0pt] \textstyle\mathcal{I}_{O}=\left\langle O,\Omega_{O},\{V_{q}^{O}:q\in\Omega_{O}\},f_{O}:O\to\prod_{q\in\Omega_{A}}V_{q}^{O}\right\rangle,\\[5.0pt] \textstyle\mathcal{I}_{E}=\left\langle E,\Omega_{E},\{V_{q}^{A}:q\in\Omega_{E}\},f_{E}:E\to\prod_{q\in\Omega_{A}}V_{q}^{E}\right\rangle.$
@@
See Figure 2 for a pictorial interpretation. A tuple in $\varphi$ is interpreted as “action $\mathtt{{Act}}_{\varphi}$ is afforded for actor $a$ described by $f_{A}(a)$ by entity $o$ described by $f_{O}(o)$ in the context $e$ described by $f_{E}(e)$”. To express this and declutter notation we agree that by actor (resp. object, environment) we always mean $\left\langle a,f_{A}(a)\right\rangle$ (resp. $\left\langle o,f_{A}(o)\right\rangle$, $\left\langle e,f_{A}(e)\right\rangle$). By an action we mean an affordance that has been labeled (given a name), and such it has been recognized as a possibility.
Definition 4.2.
A tuple $\left\langle\mathcal{I}_{A},\mathcal{I}_{O},\mathcal{I}_{E},\varphi\right\rangle$ is called an affordance structure. $\dashv$
@@
If the value is $\mathrm{Low}$ the agility is low (e.g., an actor can walk, jump to very low heights, but not run very fast); if $\mathrm{Standard}$ the actor has a <<standard>> level off agility, i.e., can run, jump (but not very high), etc.; if $\mathrm{High}$, the actor has a college varsity member agility; and if $\mathrm{Pro}$ they are on the professional level.
$V^{E}\coloneqq\{0,1\}$ where the numbers 0 and $1$ are Booleans, that is they unequivocally indicate if the environment (the playground) is adapted for one of the three activities: 0 means No, 1 means Yes. For example, for $f_{E}\colon E\to V_{\mathrm{b}}^{E}\times V_{\mathrm{h}}^{E}\times V_{\mathrm{s}}^{E}$ (where the letters ‘b’, ‘h‘ and ‘e’ abbreviate, respectively, ‘basketball’, ‘handball’ and ‘soccer’), if $f_{E}(e)=\left\langle 1,0,1\right\rangle$ it means that environment is adapted for playing basketball and soccer, but not for handball. Here, we treat adapted for playing basketball (or any of two other games) as a kind of high-level, or high-order, or aggregate property which might be described by means of simpler ones, e.g., being flat, being horizontally leveled, having at least one basketball hoop, etc. However, we simplify it as the refinement of a description of an environment (an actor, an object) seems to be dependent on the goal, or the background level of the interpreter of the components of an affordance.
We can now look at the product of the three information systems and represent an affordance by a set of vectors. In Figure 3, an affordance $\varphi$ consists of all triples of blue vectors.
Our intention is that the triples represent an affordance allowing for an action of dunking a ball. All objects in each information system are concrete entities (not equivalence classes). The actors $a_{6}$ and $a_{10}$ have the same properties but only one is in the affordance, which means that there are some hidden factors that influence the second actor, for example knee injury. $\dashv$
@@
Information system for actors
$A/_{\sim}$ $\Omega_{A}$ $p_{1}$ $p_{2}$ $p_{3}$ $p_{4}$ $p_{5}$ $p_{6}$
$a_{1}]$ $v_{a_{1}}^{p_{1}}$ $v_{a_{1}}^{p_{2}}$ $v_{a_{2}}^{p_{3}}$ $v_{a_{2}}^{p_{4}}$ $v_{a_{2}}^{p_{5}}$ $v_{a_{2}}^{p_{6}}$
$a_{2}]$ $v_{a_{2}}^{p_{1}}$ $v_{a_{2}}^{p_{2}}$ $v_{a_{2}}^{p_{3}}$ $v_{a_{2}}^{p_{4}}$ $v_{a_{2}}^{p_{5}}$ $v_{a_{2}}^{p_{6}}$
$a_{3}]$ $v_{a_{3}}^{p_{1}}$ $v_{a_{3}}^{p_{2}}$ $v_{a_{3}}^{p_{3}}$ $v_{a_{3}}^{p_{4}}$ $v_{a_{3}}^{p_{5}}$ $v_{a_{3}}^{p_{6}}$
$a_{4}]$ $v_{a_{4}}^{p_{1}}$ $v_{a_{4}}^{p_{2}}$ $v_{a_{4}}^{p_{3}}$ $v_{a_{4}}^{p_{4}}$ $v_{a_{4}}^{p_{5}}$ $v_{a_{4}}^{p_{6}}$
$a_{5}]$ $v_{a_{5}}^{p_{1}}$ $v_{a_{5}}^{p_{2}}$ $v_{a_{5}}^{p_{3}}$ $v_{a_{5}}^{p_{4}}$ $v_{a_{5}}^{p_{5}}$ $v_{a_{5}}^{p_{6}}$
$a_{6}]$ $v_{a_{6}}^{p_{1}}$ $v_{a_{6}}^{p_{2}}$ $v_{a_{6}}^{p_{3}}$ $v_{a_{6}}^{p_{4}}$ $v_{a_{6}}^{p_{5}}$ $v_{a_{6}}^{p_{6}}$
$a_{7}]$ $v_{a_{7}}^{p_{1}}$ $v_{a_{7}}^{p_{2}}$ $v_{a_{7}}^{p_{3}}$ $v_{a_{7}}^{p_{4}}$ $v_{a_{7}}^{p_{5}}$ $v_{a_{7}}^{p_{6}}$
Information system for entities
$O/_{\sim}$ $\Omega_{O}$ $q_{1}$ $q_{2}$ $q_{3}$ $q_{4}$ $q_{5}$ $q_{6}$ $q_{7}$
$o_{1}]$ $v_{o_{1}}^{q_{1}}$ $v_{o_{1}}^{q_{2}}$ $v_{o_{1}}^{q_{3}}$ $v_{o_{1}}^{q_{4}}$ $v_{o_{1}}^{q_{5}}$ $v_{o_{1}}^{q_{6}}$ $v_{o_{1}}^{q_{7}}$
$o_{2}]$ $v_{o_{2}}^{q_{1}}$ $v_{o_{2}}^{q_{2}}$ $v_{o_{2}}^{q_{3}}$ $v_{o_{2}}^{q_{4}}$ $v_{o_{2}}^{q_{5}}$ $v_{o_{2}}^{q_{6}}$ $v_{o_{2}}^{q_{7}}$
$o_{3}]$ $v_{o_{3}}^{q_{1}}$ $v_{o_{3}}^{q_{2}}$ $v_{o_{3}}^{q_{3}}$ $v_{o_{3}}^{q_{4}}$ $v_{o_{3}}^{q_{5}}$ $v_{o_{3}}^{q_{6}}$ $v_{o_{3}}^{q_{7}}$
$o_{4}]$ $v_{o_{4}}^{q_{1}}$ $v_{o_{4}}^{q_{2}}$ $v_{o_{4}}^{q_{3}}$ $v_{o_{4}}^{q_{4}}$ $v_{o_{4}}^{q_{5}}$ $v_{o_{4}}^{q_{6}}$ $v_{o_{4}}^{q_{7}}$
$o_{5}]$ $v_{o_{5}}^{q_{1}}$ $v_{o_{5}}^{q_{2}}$ $v_{o_{5}}^{q_{3}}$ $v_{o_{5}}^{q_{4}}$ $v_{o_{5}}^{q_{5}}$ $v_{o_{5}}^{q_{6}}$ $v_{o_{5}}^{q_{7}}$
$o_{6}]$ $v_{o_{6}}^{q_{1}}$ $v_{o_{6}}^{q_{2}}$ $v_{o_{6}}^{q_{3}}$ $v_{o_{6}}^{q_{4}}$ $v_{o_{6}}^{q_{5}}$ $v_{o_{6}}^{q_{6}}$ $v_{o_{6}}^{q_{7}}$
$o_{7}]$ $v_{o_{7}}^{q_{1}}$ $v_{o_{7}}^{q_{2}}$ $v_{o_{7}}^{q_{3}}$ $v_{o_{7}}^{q_{4}}$ $v_{o_{7}}^{q_{5}}$ $v_{o_{7}}^{q_{6}}$ $v_{o_{7}}^{q_{7}}$
$o_{8}]$ $v_{o_{8}}^{q_{1}}$ $v_{o_{8}}^{q_{2}}$ $v_{o_{8}}^{q_{3}}$ $v_{o_{8}}^{q_{4}}$ $v_{o_{8}}^{q_{5}}$ $v_{o_{8}}^{q_{6}}$ $v_{o_{8}}^{q_{7}}$
Information system for environments
$E/_{\sim}$ $\Omega_{E}$ $r_{1}$ $r_{2}$ $r_{3}$ $r_{4}$ $r_{5}$ $r_{6}$ $r_{7}$
$e_{1}]$ $v_{e_{1}}^{r_{1}}$ $v_{e_{1}}^{r_{2}}$ $v_{e_{1}}^{r_{3}}$ $v_{e_{1}}^{r_{4}}$ $v_{e_{1}}^{r_{5}}$ $v_{e_{1}}^{r_{6}}$ $v_{e_{1}}^{r_{7}}$
$e_{2}]$ $v_{e_{2}}^{r_{1}}$ $v_{e_{2}}^{r_{2}}$ $v_{e_{2}}^{r_{3}}$ $v_{e_{2}}^{r_{4}}$ $v_{e_{2}}^{r_{5}}$ $v_{e_{2}}^{r_{6}}$ $v_{e_{2}}^{r_{7}}$
$e_{3}]$ $v_{e_{3}}^{r_{1}}$ $v_{e_{3}}^{r_{2}}$ $v_{e_{3}}^{r_{3}}$ $v_{e_{3}}^{r_{4}}$ $v_{e_{3}}^{r_{5}}$ $v_{e_{3}}^{r_{6}}$ $v_{e_{3}}^{r_{7}}$
$e_{4}]$ $v_{e_{4}}^{r_{1}}$ $v_{e_{4}}^{r_{2}}$ $v_{e_{4}}^{r_{3}}$ $v_{e_{4}}^{r_{4}}$ $v_{e_{4}}^{r_{5}}$ $v_{e_{4}}^{r_{6}}$ $v_{e_{4}}^{r_{7}}$
$e_{5}]$ $v_{e_{5}}^{r_{1}}$ $v_{e_{5}}^{r_{2}}$ $v_{e_{5}}^{r_{3}}$ $v_{e_{5}}^{r_{4}}$ $v_{e_{5}}^{r_{5}}$ $v_{e_{5}}^{r_{6}}$ $v_{e_{5}}^{r_{7}}$
$e_{6}]$ $v_{e_{6}}^{r_{1}}$ $v_{e_{6}}^{r_{2}}$ $v_{e_{6}}^{r_{3}}$ $v_{e_{6}}^{r_{4}}$ $v_{e_{6}}^{r_{5}}$ $v_{e_{6}}^{r_{6}}$ $v_{e_{6}}^{r_{7}}$
$e_{7}]$ $v_{e_{7}}^{r_{1}}$ $v_{e_{7}}^{r_{2}}$ $v_{e_{7}}^{r_{3}}$ $v_{e_{7}}^{r_{4}}$ $v_{e_{7}}^{r_{5}}$ $v_{e_{7}}^{r_{6}}$ $v_{e_{7}}^{r_{7}}$
$e_{8}]$ $v_{e_{8}}^{r_{1}}$ $v_{e_{8}}^{r_{2}}$ $v_{e_{8}}^{r_{3}}$ $v_{e_{8}}^{r_{4}}$ $v_{e_{8}}^{r_{5}}$ $v_{e_{8}}^{r_{6}}$ $v_{e_{8}}^{r_{7}}$
$e_{9}]$ $v_{e_{9}}^{r_{1}}$ $v_{e_{9}}^{r_{2}}$ $v_{e_{9}}^{r_{3}}$ $v_{e_{9}}^{r_{4}}$ $v_{e_{9}}^{r_{5}}$ $v_{e_{9}}^{r_{6}}$ $v_{e_{9}}^{r_{7}}$
[…]
Figure 2. The triples of vectors of the same color constitute the affordance $\varphi$ and its corresponding action $\mathtt{{Act}}_{\varphi}$. We identify, respectively, actors, objects, and environments that cannot be distinguished by available properties.
Information system for actors
AA $\Omega_{A}$ height (cm) agility
$a_{1}$ 201 L
$a_{2}$ 155 P
$a_{3}$ 192 H
$a_{4}$ 190 L
$a_{5}$ 178 S
$a_{6}$ 181 P
$a_{7}$ 190 L
$a_{8}$ 201 L
$a_{9}$ 178 S
$a_{10}$ 181 P
Information system for objects
$O$ $\Omega_{O}$ diameter (in cm) weight (in kg)
$o_{1}$ 24.2 0.62
$o_{2}$ 48.2 0.75
$o_{3}$ 6.74 0.06
$o_{4}$ 20 0.41
$o_{5}$ 25.1 50
$o_{6}$ 0.4 0.02
Information system for environments
$E$ $\Omega_{E}$ basketball handball soccer
$e_{1}$ 1 1 1
$e_{2}$ 1 1 0
$e_{3}$ 1 0 1
$e_{4}$ 1 0 0
$e_{5}$ 0 1 1
$e_{6}$ 0 1 0
$e_{7}$ 0 0 1
$e_{8}$ 0 0 0
$e_{9}$ 1 1 1
$e_{10}$ 0 1 1
$\times$$\times$
Figure 3. The blue triples of vectors constitute an action $\mathtt{dunk}_{\varphi}$ of dunking a ball.

### v2

@@ 1. Introduction
The affordances of the environment are what it offers the animal, what it provides or furnishes, either for good or ill …I mean by it something that refers to both the environment and the animal in a way that no existing term does. It implies the complementarity of the animal and the environment. [14, p. 119]
@@
  • A dog fetches a stick on a meadow,
  • A basketball player dunks a ball on a playground,
  • A robot serves a meal in a restaurant,
  • A student solves a test in a classroom,
  • A mathematician writes a proof on a blackboard in their office,
@@
In this work, we intend to conceptualize affordances as ternary relations between actors, objects, and environments, taking into account their properties. [^1] Some work toward affordances as ternary phenomena (though not necessarily relations in the standard logical sense) was carried out in [21]. The main goal is to construct a framework for building logic (or logics) for reasoning with affordances. Our pre-theoretical definition of affordance is as follows:
@@
An affordance is an action possibility formed by the relationship among an actor, an object, and their environment. $\dashv$
The relevant components of our approach to this work are as follows:
  • Actors and objects have various properties, such as height, weight, color, agility, abilities, disabilities, skills, shape, and so on.
  • Environments also have properties that may be elementary (in some sense), like flat, bumpy, stifling, hot, wet, rainy, or, more complex, adapted for playing basketball, multi-level, equipped with an elevator, dangerous for health, to mention a few.
@@ 2. Information systems
If we think of a property system as describing possible states of an attribute—such as “color” or “language spoken”—we extend it by the definition of an aggregate structure: An attribute system (A-system) [23] is a structure $\mathcal{S}\coloneqq\left\langle U,\Omega,\{V_{p}:p\in\Omega\},f\right\rangle$ where
@@
    $\Omega$ is a set of property labels or attributes, and $V_{p}$ is a set of possible values of $p\in\Omega$,
@@
    $f\colon U\times\Omega\to\bigcup_{p\in\Omega}2^{V_{p}}$ is a choice function, where $f(x,p)\subseteq V_{p}$. Equivalently, we may define $f\colon U\to\prod_{p\in\Omega}2^{V_{a}}$.
So, if $p\in\Omega$ is a property label weight, then $V_{\mathit{weight}}$ may be a set of rational numbers in some interval that can serve as numerical expressions of the weight of an object (e.g., in kilograms or pounds), or any value that makes sense. It could also be some aggregated value such as “low”, “medium”, “high” etc.
We call $\left\langle U,\Omega,\{V_{p}:p\in\Omega\}\right\rangle$ the skeleton of $\mathcal{S}$. The product $U\times\prod_{p\in\Omega}2^{V_{p}}$ collects all possible vectors of value sets that can be associated with some element of $U$. An information function now picks one element from $\prod_{p\in\Omega}2^{V_{p}}$ for each $x\in U$.
Clearly, an attribute system can be regarded as a collection of the property systems. Düntsch et al., [7] introduce relational attribute system as a relational form of attribute systems, and present formalizations of various interpretations of a statement such as $x\in f(x,p)$. These are closely related to databases with incomplete information investigated by Lipski, [16]. A deduction system for reasoning in relation algebras derived from relational attribute systems was presented in [8].
An element $x\in U$ is called deterministic, if $|f(x,p)|\leq 1$ for all $p\in\Omega$ or every projection of the vector attribute to $x$ is either an empty set or a singleton subset of $V_{p}$. The set of all deterministic elements of $\mathcal{S}$ is denoted by $D_{\mathcal{S}}$. The characterization stems from the fact that the role of the choice function is to narrow down the possibilities for the values of $p$ with respect to the object $x$. If $|f(x,p)|=1$, then we know the exact value of the property $p$ for $x$, or if $|f(x,p)|=\emptyset$, then we know that $x$ does not have this property at all. This is why we call a system with $|f(x,p)|\leqslant 1$ deterministic. If $|f(x,p)|\geq 2$ the set has different possibilities of interpretation, see [7].
If $U$ is finite and $U=D_{\mathcal{S}}$, then $\mathcal{S}$ is called an information system (in the sense of Pawlak, [20]). In this case we may simplify the definition of the information function to have codomain $\bigcup_{p\in\Omega}V_{p}$. Vakarelov, [23] calls these systems “ontological information systems”, and provides logical counterparts of P-systems and A-systems based on Scott’s deduction systems. He continues to develop modal logics for similarity relations on P-systems and A-systems and shows completeness with respect to finite models, continued in [24].
@@
The properties we consider do not have to be in any way simple or elementary. We can consider complex features, such as being adapted to playing basketball. Below, we have an information system for playgrounds, each of which is characterized in terms of three complex attributes: being adapted or not (in symbols 1 or 0) to playing basketball, soccer, and volleyball.
@@
corresponds to the configuration of four environments where in $e_{1}$ one diode is active, in $e_{2}$ no diode is active, in $e_{3}$ two diodes are active, and in $e_{4}$ eight. Even in such a relatively simple environment we have
@@ 2.1. Information systems for environments
    We assume that there is one underlying environment (or context) in some pre-theoretical sense, the elements of $E$ are <<parts>> of the environment, and labels in $\Omega_{E}$ are interpreted as features of these parts. In this case, $V^{E}_{p}$ for $p\in\Omega_{E}$ is the set of values for parts of the environment.
@@
    $\Omega_{E}\coloneqq\{\mathrm{state}\}$, one label interpreted as the state of a diode.
@@
Consequently, we have sixteen different information systems based on the skeleton
$\left\langle E,\Omega_{E},\{V^{E}_{\mathrm{state}}\}\right\rangle.$
Here, there is one environment, and the elements of $E$ are parts of this environment. So this example is a representative of the perspective (E1) above. $\dashv$
@@
$\Omega_{E}\coloneqq\{\mathrm{state}_{1},\mathrm{state}_{2},\mathrm{state}_{3},\mathrm{state}_{4}\},$
@@
corresponds to the configuration of four environments where in $e_{1}$ one diode is active, in $e_{2}$ no diode is active, in $e_{3}$ two diodes are active, and in $e_{4}$ eight. Even in such a relatively simple environment we have
@@
functions $f_{E}$ and as many information systems with the skeleton $\left\langle E,\Omega_{E},\{V^{E}_{\mathrm{state}_{i}}\mid i\leqslant 4\}\right\rangle$. Observe that the skeleton from this example can be obtained from four simpler skeletons of the kind described in Example 2.4. The difference lies in the fact that in this example we can look into the structure of the elements of each $e\in E$. $\dashv$
@@ 3. Rough sets
Given an information system $\mathcal{I}\coloneqq\left\langle U,\Omega,\{V_{p}\mid p\in\Omega\},f\right\rangle$, for all $a,b\in U$
$a\mathrel{\theta}b\;\mathrel{\mathord{:}\mathord{\Longleftrightarrow}}\;f(a)=f(b),$
that is, $a$ and $b$ have the same value vector. In this case, we say that $a$ is similar to $b$ (or that $a$ and $b$ are similar). Obviously, in the case of the deterministic systems, $a$ and $b$ are similar if they have precisely the same values for all the property labels in $\Omega$. $\dashv$
Clearly, $\theta$ is an equivalence relation on $U$ whose classes we denote by $a]_{\theta}$ or simply by $[a]$ if no confusion can arise. By means of $\theta$, we can make use of a concept of a rough set to incorporate the granularity of the world into our formalism. Given a subset $X$ of $U$ and an element $a\in U$, when can we say that certainly $a\in X$ or certainly $a\not\in X$? Since we know the world only up to the classes of $\theta$ we can say that
@@
$\overline{X}=\{a_{1},a_{3},a_{4},a_{5},a_{7},a_{8},a_{9}\}\quad\text{and}\quad\underline{X}=\{a_{3},a_{5},a_{9}\}.$
Therefore, the boundary of $X$ is the set $\{a_{1},a_{4},a_{7},a_{8}\}$.
Next, consider $X^{\prime}\coloneqq\ X\setminus\{a_{3},a_{5}\}=\{a_{1},a_{4},a_{9}\}$. Then, $\underline{X^{\prime}}=\emptyset$ and we cannot say with certainty that $a\in X^{\prime}$ for any $a\in A$. Dually, if $X^{\prime\prime}=A\setminus X^{\prime}=\{a_{2},a_{3},a_{5},a_{6},a_{7},a_{8}\}$, then $\overline{X^{\prime\prime}}=A$, and any element of $A$ is possibly in $X^{\prime\prime}$ with the information at hand. $\dashv$
The previous example shows that approximation operators can yield counterintuitive results. If, say, $\theta$ is an equivalence relation on $A$ with two very large classes $X$ and $Y$, and $x\in X,y\in Y$, then $\underline{A\setminus\{x,y\}}=\emptyset$, and $\overline{\{x,y\}}=A$. To handle such situations, in particular, when rough set methods are used for classification or supervised learning, additional statistical methods are required, such as approximation quality, significance testing, or information theoretic entropy. An overview of such methods can be found in [13].
@@ 4. Operationalizing affordances
The basic setup of an affordance relation consists of a set $U$ of an agent’s abilities, a set $E$ of features of the environment, and a binary relation $R\subseteq U\times E$. Chemero, [2, p. 189] writes
@@
Above, $\mathtt{{Act}}_{\varphi}$ is thought of as an action related to the triple $\varphi$ (further below, we give a more formal explanation).
@@
$\mathcal{I}_{A}=\left\langle A,\Omega_{A},\{V_{p}^{A}:p\in\Omega_{A}\},f_{A}:A\to\prod_{p\in\Omega_{A}}V_{p}^{A}\right\rangle,\\[5.0pt] \textstyle\mathcal{I}_{O}=\left\langle O,\Omega_{O},\{V_{p}^{O}:p\in\Omega_{O}\},f_{O}:O\to\prod_{p\in\Omega_{A}}V_{p}^{O}\right\rangle,\\[5.0pt] \textstyle\mathcal{I}_{E}=\left\langle E,\Omega_{E},\{V_{p}^{A}:p\in\Omega_{E}\},f_{E}:E\to\prod_{p\in\Omega_{A}}V_{p}^{E}\right\rangle.$
@@
See Figure LABEL:fig:afford for a pictorial interpretation. A tuple in $\varphi$ is interpreted as “action $\mathtt{{Act}}_{\varphi}$ is afforded for actor $a$ described by $f_{A}(a)$ by entity $o$ described by $f_{O}(o)$ in the context $e$ described by $f_{E}(e)$”. To express this and declutter notation we agree that by actor (resp. object, environment) we always mean $\left\langle a,f_{A}(a)\right\rangle$ (resp. $\left\langle o,f_{O}(o)\right\rangle$, $\left\langle e,f_{E}(e)\right\rangle$). By an action we mean an affordance that has been labeled (given a name), and such it has been recognized as a possibility.
@@
If the value is $\mathrm{Low}$ the agility is low (e.g., an actor can walk, jump to very low heights, but not run very fast); if $\mathrm{Standard}$ the actor has a <<standard>> level of agility, i.e., can run, jump (but not very high), etc.; if $\mathrm{High}$, the actor has a college varsity member agility; and if $\mathrm{Pro}$ they are on the professional level.
$V^{E}\coloneqq\{0,1\}$ where the numbers 0 and $1$ are Booleans, that is they unequivocally indicate if the environment (the playground) is adapted for one of the three activities: 0 means No, 1 means Yes. For example, for $f_{E}\colon E\to V_{\mathrm{b}}^{E}\times V_{\mathrm{h}}^{E}\times V_{\mathrm{s}}^{E}$ (where the letters ‘b’, ‘h‘ and ‘s’ abbreviate, respectively, ‘basketball’, ‘handball’ and ‘soccer’), if $f_{E}(e)=\left\langle 1,0,1\right\rangle$ it means that environment is adapted for playing basketball and soccer, but not for handball. Here, we treat adapted for playing basketball (or any of two other games) as a kind of high-level, or high-order, or aggregate property which might be described by means of simpler ones, e.g., being flat, being horizontally leveled, having at least one basketball hoop, etc. However, we simplify it as the refinement of a description of an environment (an actor, an object) seems to be dependent on the goal, or the background level of the interpreter of the components of an affordance.
We can now look at the product of the three information systems and represent an affordance by a set of vectors. In Figure LABEL:fig:dunk, an affordance $\varphi$ consists of all triples of blue vectors.
Our intention is that the triples represent an affordance allowing for an action of dunking a ball. All objects in each information system are concrete entities (not equivalence classes). The actors $a_{6}$ and $a_{10}$ have the same properties but only one is in the affordance, which means that there are some hidden factors that influence the second actor, for example, knee injury. $\dashv$